OpenStudy (anonymous):

Given the following endpoints (A & B) and midpoint (M) on a number line, find the coordinates for A & B. M(-11) A(2x + 3) B(5x + 3)

OpenStudy (amistre64):

m-a = m+b -a = b

OpenStudy (anonymous):

?

OpenStudy (amistre64):

yeah, thats what i was thinking too lol

OpenStudy (amistre64):

the distance from m to a is equal to m to b

OpenStudy (amistre64):

when m = 0 we got a = b then

OpenStudy (amistre64):

2x+3 = 5x+3 ....maybe 0 = 3x x = 0

OpenStudy (amistre64):

-11 +3 = -8 -11 - 3 = -14

OpenStudy (amistre64):

it makes better sense to me now :) i really shouldnt doubt myself like that in front of my other personalities. sends a bad message :)

OpenStudy (anonymous):

that's isnt an answer choice

OpenStudy (anonymous):

okay, all you have to do is these: First, the point 2x+3 and 5x+3 has a midpoint of -11. (2x+3+5x+3)/2=-11 (2x+3+5x+3)=-22 7x+6=-22 7x=-22-6 7x=-28 x=-4 A=2x+3 =2(-4)+3 =-8+3 =-5 B=5x+3 =5(-4)+3 =-20+3 =-17